OpenStudy (anonymous):

What is the sum of a 7–term geometric series if the first term is –11, the last term is –171,875, and the common ratio is –5?

OpenStudy (anonymous):

-143231

OpenStudy (anonymous):

sum = -11 * (-5)^7 - 1 --------- -5-1

OpenStudy (anonymous):

@jimmyrep : \( 1 -(-5) \) isn't ?

OpenStudy (anonymous):

r is less than 1

OpenStudy (anonymous):

yea - u r right it should be 1 - r since r < 1

OpenStudy (anonymous):

sum = -11 * (1 - (-5)^7) -------- 1 - (-5)

OpenStudy (anonymous):

= -143231

OpenStudy (anonymous):

thank you! so you do the fraction up top first then divide it by the bottom and then multiply right orr?

OpenStudy (anonymous):

yea - you can do the calculation that way you know something though- although its usual to use 1-r^n and 1-r if r < 1 i'm not sure that it matters - i think it will give same answer - i'll check that out with this question

OpenStudy (anonymous):

yea - it gives same answer

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

@jimmyrep: Try taking limit when \( n \to \infty \)

OpenStudy (anonymous):

yes - i know what you are saying - its only that it works in a case like this one - however as a general rule you should use 1 - r when r < 1

OpenStudy (anonymous):

:)